# Multi-stream Physics Hybrid Networks for solving Navier-Stokes equations

### v1

@@ I Introduction
Machine learning can be greatly improved by using quantum technologies. In gaitan2020finding quantum computing is used within a similar problem. The performance of existing machine learning models is limited by high computational resource requirements. Quantum computing can improve the learning process of classical models, allowing for better accuracy in predicting the target function with fewer iterations dunjko2018machine ; qml_review_2023 ; Neven2012QBoostLS ; rebentrost2014quantum ; saggio2021experimental ; kordzanganeh2023parallel ; kordzanganeh2023benchmarking . In many industrial and scientific fields, such as pharmaceutical sagingalieva2023hybrid ; gircha2021training , aerospace rainjonneau2023quantum , automotive sagingalieva2023hyperparameter , and even financial alcazar2020classical ; coyle2021quantum ; pistoia2021quantum ; emmanoulopoulos2022quantum ; raj2023quantum quantum technologies can provide significant improvements to existing classical methods. Many traditional machine learning tasks such as image processing senokosov2023quantum ; li2022image ; naumov2025tensor ; riaz2023accurate ; lusnig2024hybrid and natural language processing hong2022qspeech ; lorenz2023qnlp ; coecke2020foundations ; meichanetzidis2020grammar have already demonstrated the broad application prospects of quantum methods. The goal of this work is to explore the feasibility of applying quantum machine learning to the new and emerging field of physical modeling using neural networks.
@@ II Kovasznay flow problem formulation
We will be solving the Navier-Stokes equations in rectangular domain $\Omega=[-0.5,1.0]\times[-0.5,1.5]$. Kovasznay flow model describes the fluid flow behind two dimensional grid. The flow is laminar and is governed by $2$D Navier-Stokes equations:
@@ IV.1 Multi-stream Hybrid Network
[…]
@@
The classical layer is a small fully connected neural network with $h=10$ neurons in the hidden layer, with activation functions ReLU standing between the layers. The quantum layer is a parameterized quantum two-qubit circuit, shown in Fig. 1. The use of parameterized quantum circuits is widespread in quantum machine learning because it is a synthesis of the best ideas of classical machine learning and quantum computing zhao2019qdnn ; dou2021unsupervised ; sagingalieva2023hybrid ; kordzanganeh2023exponentially .
@@
The chosen quantum circuit is quite simple, but demonstrates high efficiency. Training of quantum circuits on a simulator (classical computer) usually takes quite a long time as it is impossible to use backpropagation rumelhart1986learning when calculating gradients, so simple quantum circuits are trained faster. Current NISQ (noisy intermediate-scale quantum) devices, i.e., real quantum processors used today, are also not yet capable of high-precision computations on deep quantum circuits, so simplicity is a big advantage for us. Moreover, using a large number of qubits or a large depth in variational quantum circuits often leads to damped gradients – Barren plateaus mcclean2018barren ; cerezo2021cost .
@@
$\begin{split}\mathcal{L}_{v_{x}}&=\frac{1}{n}\sum_{i=1}^{n}(v_{x}^{\text{p}}(\bm{r}_{i})-v_{x}^{\text{e}}(\bm{r}_{i}))^{2},\\ \mathcal{L}_{v_{y}}&=\frac{1}{n}\sum_{i=1}^{n}(v_{y}^{\text{p}}(\bm{r}_{i})-v_{y}^{\text{e}}(\bm{r}_{i}))^{2},\\ \mathcal{L}_{p}&=\frac{1}{n}\sum_{i=1}^{n}(p^{\text{p}}(\bm{r}_{i})-p^{\text{e}}(\bm{r}_{i}))^{2},\\ \end{split}$ (13)
@@ IV.3 Training
Figure 2: Training results of multi-stream network models after 100 epochs in data-driven learning. Here Exact, MCN and MHN denote the exact solution, classical multi-stream and hybrid multi-stream networks respectively, and $v_{x},v_{y},p$ are velocity and pressure projections. The classical MCN model does well in predicting the $p$ function, but cannot approximate the periodic functions $v_{x}$ and $v_{y}$. Hybrid MHN model approximates all solution functions equally well, with even fewer trainable parameters.

### v3

@@ I Introduction
Machine learning can be greatly improved by using quantum technologies. In gaitan2020finding quantum computing is used within a similar problem. The performance of existing machine learning models is limited by high computational resource requirements. Quantum computing can improve the learning process of classical models, allowing for better accuracy in predicting the target function with fewer iterations dunjko2018machine ; qml_review_2023 ; Neven2012QBoostLS ; rebentrost2014quantum ; saggio2021experimental ; kordzanganeh2023parallel ; kordzanganeh2023benchmarking . In many industrial and scientific fields, such as pharmaceutical sagingalieva2023hybrid ; gircha2021training , aerospace rainjonneau2023quantum , automotive sagingalieva2023hyperparameter , and financial alcazar2020classical ; coyle2021quantum ; pistoia2021quantum ; emmanoulopoulos2022quantum ; raj2023quantum quantum technologies can provide improvements to existing classical methods. Many traditional machine learning tasks such as image processing senokosov2023quantum ; li2022image ; naumov2025tensor ; lusnig2024hybrid ; riaz2023accurate ; sagingalieva2025hybrid , time series forecasting kurkin2025forecasting ; sagingalieva2025photovoltaic ; lee2025predictive , and natural language processing hong2022qspeech ; lorenz2023qnlp ; coecke2020foundations ; meichanetzidis2020grammar have already demonstrated the broad application prospects of quantum methods. The goal of this work is to explore the feasibility of applying quantum machine learning to the new and emerging field of physical modeling using neural networks.
@@ II Kovasznay flow problem formulation
We will be solving the Navier-Stokes equations in rectangular domain $\Omega=[-0.5,1.0]\times[-0.5,1.5]$. Kovasznay flow model describes the fluid flow behind two dimensional grid. The flow is laminar and is governed by 22D Navier-Stokes equations:
@@ IV.1 Multi-stream Hybrid Network
Here we introduce MHN – Multi-stream Hybrid Network (without the “physics” part). MHN model is divided into 3 independent parts (PHNs), respectively for predicting $u,v$ and $p$, refer to Fig. 1. This is motivated by the fact that these functions describe different physical values and their scale can differ. Each of PHN layers accept the $(x,y)$ coordinates. Then the outputs of these layers, respectively “Quantum Output” (Qout) and “Classical Output” (Cout) are affinely transformed, with an addition of cross-term $\texttt{Qout}\cdot\texttt{Cout}$
@@
The classical layer is a small fully connected neural network with $h=10$ neurons in the hidden layer, with activation functions ReLU standing between the layers. The quantum layer is a parameterized quantum two-qubit circuit, shown in Fig. 1. The use of parameterized quantum circuits is widespread in quantum machine learning because it is a synthesis of the best ideas of classical machine learning and quantum computing zhao2019qdnn ; sagingalieva2023hybrid ; dou2021unsupervised ; kordzanganeh2023exponentially ; haboury2024information ; patapovich2025superposed .
@@
The chosen quantum circuit is quite simple, but demonstrates high efficiency. Training of quantum circuits on a simulator kuzmin2025tqml (classical computer) usually takes quite a long time as it is impossible to use backpropagation rumelhart1986learning when calculating gradients, so simple quantum circuits are trained faster. Current NISQ (noisy intermediate-scale quantum) devices, i.e., real quantum processors used today, are also not yet capable of high-precision computations on deep quantum circuits, so simplicity is a big advantage for us. Moreover, using a large number of qubits or a large depth in variational quantum circuits often leads to damped gradients – Barren plateaus mcclean2018barren ; cerezo2021cost .
@@
[…]
@@ IV.3 Training
Figure 2: Training results of multi-stream network models after 100 epochs in data-driven learning. Here Exact, MCN and MHN denote the exact solution, classical multi-stream and hybrid multi-stream networks respectively, and $v_{x},v_{y},p$ are velocity and pressure projections. The MCN model does well in predicting the $p$ function, but cannot approximate the periodic functions $v_{x}$ and $v_{y}$. The MHN model approximates all solution functions equally well, with even fewer trainable parameters.